A car and a bus set out at 6 P.M. from the same point headed in the same direction. The average speed of the car is twice the average speed of the bus. In 2 h, the car is 68 mi ahead of the bus. Find the rate of the car.
Answer by Theo(13342)

r * t = d
r = rate
t = time
d = distance.

if r = rate of the bus, then 2r = rate of the car.
if x is the distance that the bus travels, then x + 68 is the distance the car traveled.

when time = 2, you get:

2r * 2 = x + 68 for the car.
r * 2 = x for the bus.

simplify to get:

4r = x + 68 for the car.
2r = x for the bus.

subtract the second equation from the first to get:

2r = 68

solve for r to get:

r = 35

the rate of the bus = 34 miles per hour.
the rate of the car = 68 miles per hour.

in 2 hours, the car travels 68 * 2 = 136 miles.
in 2 hours, the bus travels 34 * 2 = 68 miles.

the car travels 68 miles more than the bus in the same time.

your solution is the rate of the car = 68 miles per hour.